A HIGH-RESOLUTION STUDY OF THE FOUR LOWEST FUNDAMENTAL BANDS AND ACCURATE DETERMINATION OF THE GROUND STATE CONSTANTS OF H_{3}Si^{37}Cl

\underline{G. GRANER}, S. BOSC, Y. HENNEQUIN, Laboratoire de Physique Moléculaire et Applications, CNRS, Bât. 350, Campus d’Orsay, F-91405 Orsay Cédex, France; H. BÜRGER, and S. BAILLEUX, Anorganische Chemie, FB 9, Universität-Gesamthochschule, D-42097 Wuppertal, Germany.

In order to determine the structure and force field of silyl chloride, the infrared spectra of several isotopomers were studied. After monoisotopic D_{3}Si^{35}Cl (previous communication), we report here on the analysis of the monoisotopic H_{3}Si^{37}Cl species. Here too, the two fundamental bands ν_{3} (543.968cm^{−1}) and ν_{6} (663.736 cm^{−1}) are linked by a Coriolis resonance but its effects are less important than for D_{3}SiCl and especially than for the two nearly degenerate fundamental bands ν_{2} (947.982 cm^{−1}) and ν_{5} (950.657 cm^{−1}). The ’normal’ ground state constants B_{0}, D_{J}^{0}, D_{JK}^{0}, H_{J}^{0}, H_{JK}^{0}, and D_{KJ}^{0} were deduced from more than 6000 GSCD. As for the constants A_{0} and D_{K}^{0}, they were obtained by the same method as for D_{3}Si^{35}Cl. Calibration errors were detected and corrected by checking the closed loop between ν_{6}, (2ν_{6})^{0} and (2ν_{6})^{0} - ν_{6}. The values obtained are A_{0} = 2.84564 cm^{−1} and D_{K}^{0} = 2.54× 10^{−6} cm^{−1} (provisional values). Upper state energies were also determined for ν_{3}, ν_{6}, ν_{2}, ν_{5}, (2ν_{6})^{±2}, (2ν_{6})^{0}, 2ν_{3}, and ν_{3} + ν_{6}. Some interesting features of the hot bands ν_{3} + ν_{6} - ν_{3} and ν_{3} + ν_{6} - ν_{6} will also be reported.